Two moles of an ideal gas are allowed to expand reversibly and isothermally at 300 K from a
pressure of 1 atm to a pressure of 0.1 atm. What is the change in Gibbs free energy?�
1
Expert's answer
2021-08-10T05:03:55-0400

Solution:

Consider the isothermal expansion of an ideal gas: dG = VdP

From the ideal gas law, V = nRT/P

So, dG = (nRT/P)dP


Integrating both sides,



ΔG = (2 mol) × (8.314 J mol-1 K-1) × (300 K) × ln(0.1 atm / 1 atm) = -11486.2 J = -11.49 kJ

ΔG = -11.49 kJ

Answer: The change in Gibbs free energy (ΔG) is -11.49 kJ